# Densities Of Primes And Primitive Roots

### v13

@@ 5.2 FFT of Power Exponential Sum with Relatively Prime Index
where $\delta>0$ is a small number and the implied constant is independent of $b\in[1,p-1]$.
@@
confer equations (5.2), (5.3) and (5.1) for more details. In particular, for $b=1$,
@@
$\hskip 180.67499pt\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right).$
@@
$\Bigg|1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi a\tau^{s}}{p}}\Bigg|=\Bigg|1-\sum_{1\leq s\leq p-1}\chi(s)\psi(s)\Bigg|$ $\leq$ $2p^{1/2}\log p,$ (5.15)
@@
where the implied constant depends on the number of divisors of p1p-1. ∎
@@
where $\delta>0$ is a small number and the implied constant is independent of $b\in[1,p-1]$.
@@ 6 Evaluation of the Main Term
Finite sums and products over the primes numbers occur on various problems concerned with primitive roots. These sums and products often involve the normalized totient function $\varphi(n)/n=\prod_{p|n}(1-1/p)$ and the corresponding estimates, and the asymptotic formulas.
@@
([45, Lemma 1]) Let $x\geq 1$ be a large number, and let $\varphi(n)$ be the Euler totient function. Then
$\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}=a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$ (6.1)
where $\operatorname{li}(x)$ is the logarithm integral, $a_{1}=0.373955\ldots$, and $B>1$ is an arbitrary constant, as $x\rightarrow\infty$.
A more general version of this Lemma is proved in [47], and related discussions are given in [32, p. 16]. The generalization to number fields appears in [26]. These results are ubiquitous in various results in Number Theory. Here, the logarithm integral is defined by
$\operatorname{li}(x)=\int_{2}^{x}\frac{1}{\log t}dt=\frac{x}{\log x}+O\left(\frac{x}{\log^{2}x}\right).$ (6.2)
@@
coincides to the average density of primitive roots modulo $p$. The average density of primitive roots was proved in [20] and [45], (the heuristic is due to Artin, and the average density is known as Artin constant).
@@
$\sum_{p\leq x}\frac{1}{p}\sum_{\gcd(n,p-1)=1}1=a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right).$ (6.4)
@@
A routine rearrangement gives
$\sum_{p\leq x}\frac{1}{p}\sum_{\gcd(n,p-1)=1}1$ $=$ $\sum_{p\leq x}\frac{\varphi(p-1)}{p}$
$=$ $\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}.$
To proceed, apply Lemma 6.1 to reach
$\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}$ $=$ $a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right)-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}$ (6.6)
$=$ $a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$
where the second finite sum
$\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}=O\left(\log\log x\right)$ (6.7)
is absorbed into the error term, which has $B>1$ as an arbitrary constant. ∎
@@ 7 Estimate for the Error Term
The upper bounds for exponential sums over subsets of elements in finite rings $\mathbb{Z}/m\mathbb{Z}$ or finite fields $\mathbb{F}_{p}$ studied in Section 5 are used to estimate the error term $E(x)$ in the proof of Theorem 1.1.
@@
8 Densities of Primes and Primitive Roots
The representations of the characteristic function of primitive roots, Lemma 3.2, and Lemma 3.3, easily detect certain local and global properties of the elements $u\in\mathbb{F}_{p}$ in a finite field. Exempli gratia, it vanishes
[…]
Ergo, the constraints $u\neq\pm 1,v^{2}$ with $v\in\mathbb{Z}$, are necessary global constraints to be a primitive element in $\mathbb{F}_{p}$, $p\geq 2$ an arbitrary prime. But the constraints $u\neq v^{q},q\in\mathcal{Q}$, where $\mathcal{Q}$ is a finite subset of primes, are not necessary global constraints since there are infinitely many primes for which $q\nmid p-1,q\in\mathcal{Q}$. The requirement of being dth power nonresidues $\bmod p$ for all $d\mid p-1$, which is equivalent to the definition of primitive root, are necessary local properties.
@@ 8 Densities of Primes and Primitive Roots
A formula for computing the density constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is derived in [25, p. 220]. This formula specifies the density as follows. Let $u=(st^{2})^{k}\neq\pm 1,v^{2}$ with $s$ squarefree, and $k\geq 1$, and let
@@
In the case of a decimal base $u=10$, the density (6.3), which is computed in [49], is
@@
Information on the determination of the constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is provided in the next section and in [25, p. 220].
@@ 9 Application To Repeated Decimals
are also important in the calculation of the period modulo $p^{2}$, confer [41, p. 333], [DK11] for other details.
@@
10 Alternative Estimate for the Error Term
The altrrnative estimate for the error term $E(x)$ ocurring in the proof Theorem 1.1 is a recently new proof. It is significantly shorter than the original proof in Section 7.
Lemma 10.1.
$\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\begin{subarray}{c}1\leq s\leq p-1\\ \gcd(s,p-1)=1\end{subarray}}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}=O\Bigg(xe^{-c_{4}\sqrt{\log x}}\Bigg),$ (10.1)
where $c>0$ is an absolute constant, for all sufficiently large numbers $x\geq 1$.
Proof.
Employing the characteristic function
[…]
of relatively prime integers to remove the double index in the first line of (10) and switching the order of summation in the second line yield
$E(x)$ $=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\begin{subarray}{c}1\leq s\leq p-1\\ \gcd(s,p-1)=1\end{subarray}}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{1\leq s\leq p-1,}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}\sum_{\begin{subarray}{c}d\mid s\\ d\mid p-1\end{subarray}}\mu(d)$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\sum_{\begin{subarray}{c}1\leq s\leq p-1,\\ d\mid s\end{subarray}}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\sum_{1\leq m\leq(p-1)/d,}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}.$
By hypothesis $\tau^{s}-u=\tau^{dm}-u\neq 0$ for any $s\geq 1$ such that $\gcd(s,p-1)=1$. Thus, evaluating the inner sum indexed by $t$ followed by the evaluation of the resulting inner sum indexed by $m$ yield
$E(x)$ $=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\sum_{1\leq m\leq(p-1)/d}(-1)$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\left[\frac{p-1}{d}\right]$
$=$ $\sum_{x\leq p\leq 2x}\frac{p-1}{p}\sum_{d\leq p-1}\frac{\mu(d)}{d}-\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\left\{\frac{p-1}{d}\right\}\,$
where the last line follows from the largest integer function $z]=z-\{z\}$ identity.
Applying Theorem 11.1 to the first double finite sum and Lemma 11.1 to the second double finite sum return
$E(x)$ $=$ $O\Bigg(\sum_{x\leq p\leq 2x}e^{-c_{1}\sqrt{\log p}}\Bigg)+\sum_{x\leq p\leq 2x}\frac{1}{p}\Bigg(-1+O\bigg(pe^{-c_{2}\sqrt{\log p}}\bigg)\Bigg)$ (10.7)
$=$ $O\Bigg(\sum_{x\leq p\leq 2x}e^{-c_{3}\sqrt{\log p}}\Bigg)+O\bigg(\frac{1}{\log x}\bigg)$
$=$ $O\Bigg(xe^{-c_{4}\sqrt{\log x}}\Bigg),$
where the last line follows from the prime number theorem on the short interval $[x,2x]$ and $c_{1},\ldots,c_{4}>0$ are constants. ∎
11 Average Orders of Mobius Functions
Theorem 11.1.
If $\mu:\mathbb{N}\longrightarrow\{-1,0,1\}$ is the Mobius function, then, for any large number $x\geq 1$, the following statements are true.
  1. (i)
    $\sum_{n<x}\mu(n)=O\left(xe^{-c\sqrt{\log x}}\right)$,
    unconditionally,
  2. (ii)
    $\sum_{n<x}\frac{\mu(n)}{n}=O\left(e^{-c\sqrt{\log x}}\right),$
    unconditionally,
where $c>0$ is an absolute constant.
Proof.
See [12, Equation 27.11.12], [11, p. 6], [36, p. 182], et alii. ∎
Lemma 11.1.
If $x\geq 1$ is a large number, $\mu$ is the Mobius function and $\{z\}$ is the fractional part function, then
$\sum_{n\leq x}\mu(n)\bigg\{\frac{x}{n}\bigg\}=-1+O\left(xe^{-c\sqrt{\log x}}\right),$ (11.1)
where $c>0$ is an absolute constant.
Proof.
Start with the identity
$1$ $=\sum_{n\leq x}\mu(n)\bigg[\frac{x}{n}\bigg]$ (11.2)
$=x\sum_{n\leq x}\frac{\mu(n)}{n}-\sum_{n\leq x}\mu(n)\bigg\{\frac{x}{n}\bigg\},$
see [2, Theorem 3.12]. Applying Theorem 11.1 to the first finite sum on the left side and rearranging it complete the verification. ∎

### v14

@@ 4.1 Finite Fourier Transform of the Power Exponential Sum
where $\delta>0$ is a small number and the implied constant is independent of $a,b\in[1,p-1]$.
@@
confer equations (4.2), (4.3) and (4) for more details. In particular, for $b=1$,
@@
$\hskip 180.67499pt\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)$
$=$ $-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}+\widehat{R(a)}$.
@@
$\Bigg|1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\Bigg|=\Bigg|1-\sum_{1\leq s\leq p-1}\chi(s)\psi(s)\Bigg|$ $\leq$ $2p^{1/2}\log p,$ (4.14)
@@
where the implied constant depends on the number of divisors of p1p-1 and it is independent of $a,b\in[1,p-1]$. ∎
@@
where $\delta>0$ is a small number and the implied constant is independent of $a,b\in[1,p-1]$.
@@ 5 Evaluation of the Main Term
The asymptotic formula for the main term occurring in the proof of Theorem 1.1 is evaluated in this section.
@@
([42, Lemma 1]) Let $x\geq 1$ be a large number, and let $\varphi(n)$ be the Euler totient function. Then
$\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}=a_{1}\operatorname{li}(x)+O\left(\frac{x}{(\log x)^{c}}\right),$ (5.1)
where $\operatorname{li}(x)$ is the logarithm integral, $a_{1}=0.373955\ldots$, and $c>1$ is an arbitrary constant, as $x\rightarrow\infty$.
A more general version of this Lemma is proved in [45], and related discussions are given in [29, p. 16]. The generalization to number fields appears in [23]. These results are ubiquitous in various results in Number Theory. Here, the logarithm integral is defined by
$\operatorname{li}(x)=\int_{2}^{x}\frac{1}{\log t}dt=\frac{x}{\log x}+O\left(\frac{x}{(\log x)^{2}}\right).$ (5.2)
@@
coincides to the average density of primitive roots modulo $p$. The average density of primitive roots was proved in [42], (the heuristic is due to Artin, and the average density is known as Artin constant).
@@
$\sum_{p\leq x}\frac{1}{p}\sum_{\gcd(n,p-1)=1}1=a_{1}\operatorname{li}(x)+O\left(\frac{x}{(\log x)^{c}}\right),$ (5.4)
where $c>1$ is an arbitrary constant, as $x\rightarrow\infty$.
@@
A routine rearrangement and an application of Lemma 5.1 give
$\sum_{p\leq x}\frac{1}{p}\sum_{\gcd(n,p-1)=1}1$ $=$ $\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}$
$=$ $a_{1}\operatorname{li}(x)+O\left(\frac{x}{(\log x)^{c}}\right),$
where $c>1$ is an arbitrary constant. ∎
@@ 6 Estimate for the Error Term
The upper bounds for exponential sums over subsets of elements in finite fields $\mathbb{F}_{p}$ studied in Section 4 are used to estimate the error term $E(x)$ in the proof of Theorem 1.1.
@@
7 Proof of Main Result
The counting function for the number of primes $p\in[x,2x]$, with a fixed primitive $u=\pm 1,v^{2}$, is defined by
$\pi(u,x)=\#\{x\leq p\leq 2x:\operatorname{ord}_{p}u=p-1\}.$ (7.1)
@@ 7 Proof of Main Result
A formula for computing the density constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is derived in [21, p. 220]. This formula specifies the density as follows. Let $u=(st^{2})^{k}\neq\pm 1,v^{2}$ with $s$ squarefree, and $k\geq 1$, and let
@@
In the case of a decimal base $u=10$, the average density (1.1), which is computed in [46], is
@@
Information on the determination of the constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is provided in the next section and in [21, p. 220]. The range of values of the density
$\frac{84}{85}a_{1}\leq\delta(u)\leq\frac{6}{5}a_{1}$ (7.9)
is studied in [34], where the constant $a_{1}$ is defined in (5.3).
@@ 8 Application To Repeated Decimals
are also important in the calculation of the period modulo $p^{2}$, confer [40, p. 333], [10] for other details.